What is the probability of a coin tossed 3 times to get 3 heads?
If you look at this in a "fair way," it would be 1/8. However, realistically, it has been proven that the heads side weighs slightly more that the tails side so that percentage (12.5%) would decrease.
What are odds of rolling 14 with 4 dice?
ways of getting 14
1,1,6,6 6 arrangements (1,6,1,6; 1,6,6,1; 1,1,6,6; 6,6,1,1; 6,1,6,1; 6,1,1,6)
1,2,5,6 24 arr.
1,3,4,6 24 arr.
1,3,5,5 12 arr.
1,4,4,5 12 arr.
2,2,4,6 12 arr.
2,2,5,5 6 arr.
2,3,3,6 12 arr.
2,3,4,5 24 arr.
2,4,4,4 4 arr.
3,3,3,5 4 arr.
3,3,4,4 6 arr.
Total: 146
146*(1/6)4 = 0.1127... so slightly over 1/9
What is the probability of drawing a red king from a deck of 52 playing cards?
50/2
Since: there are only 2 red kings (hearts and diamonds) and there are 50 that ARE NOT red kings.
Is the standard deviation the same as the mean of absolute distances from the mean?
no the standard deviation is not equal to mean of absolute distance from the mean
What is a good sentence for probability?
There is a high probability that this sentence will not qualify as a good sentence.
The probability of you understanding this answer is slim... does that help?? That is how you use it in a sentance.
** Jeez, that's mean, and OP, check out people who have the same exact question as you on here.. "What is probability" rather than how.
Undefined; large
What is a sample space diagram?
A sample space diagram is usually a Venn diagram with the event(s). See the related link.
It is 95%
What is the probability that it will go dark tonight?
It depends on where you are. During the summer in the Northern hemisphere, if you are North of the Arctic circle, the answer is 0.
How is paired t-test different from a general t-test with two samples?
In a general t-test, there is no relationship between the members of one sample and the other. In a paired t-test they are connected in some way so that they are likely to give similar outcomes. This means that more of the difference between them can be attributed to the "treatment".
How do you calculate Q1 Q3 IQR and the standard deviation?
Suppose there are n observations. Put them in ascending order (smallest first) of size.
Calculate k = n/4. Round up to the next integer, if necessary.
Then Q1 is the kth observation in the ordered sets.
Also Q3 is the 3kth observation in the ordered sets.
IQR = Q3 - Q1
Calculation of the standards deviation is a lot more work.
First find the mean = sum of all the observations, divided by the number of observations. Call that number M.
Next find the mean "sum of squares", MSS. Square the value of each observation and add them together. Then divide this sum by the number of observations.
Then the Variance is V = MSS - M2
Finally, the standard deviation is sqrt(V).
The variance of the estimate for the mean would be reduced.
What is the purpose of studying probability and Queueing theory?
An obvious answer is to allow an organisation to decide how many people they should employ for serving customers. That may be tellers in a bank or post office, or telephone operators at a call centre.
Obviously, employing too many is a waste of the organisations resource. Unfortunately, most organisations do not realise that not employing enough, so that the customers are left waiting (or left on hold and forced to listen to some revolting "music") does not do much for their image. But hey, they saved some money, so they're good!
What is probability of an impossible event?
If the event is truly impossible, it has a zero probability of occurring. The definition of impossible is "not able to occur, exist, or be done," so that means there's no chance of it happening.
What is the probability that a white sock would be pulled out a drawer of all white socks?
Whenever there is only one possible outcome of an experiment, the probability of this outcome is 1, which you can also write as 100%.
How many queen of diamond given face card?
In an ordinary deck of cards there is only one queen of diamonds.
How to decrease margin of error in probability sampling?
With probability sampling you have no control over the units that are sampled. So the only way to reduce the margin of error is to increase the size of the sample.
How do you respond if partner responds 2 Hearts to your 1Spade opening bid?
It depends on what card game you are playing that will affect how cards are played. Playing 2 Hearts to your 1Spade opening bid would be difficult because there is no 1 Spade.
How many different four person committees can be formed from a group of 9 people?
(9 x 8 x 7 x 6)/(4 x 3 x 2 x 1) = 126committees.
What is the probability of flipping 4 times and landing on heads each time?
The probability of heads is 0.5 each time.
The probability of four times is (0.5 x 0.5 x 0.5 x 0.5) = 0.0625 = 1/16 = 6.25% .